OpenStudy (anonymous):

Simplify Squareroot of 20

OpenStudy (anonymous):

Example: simplify √18 √18 = √(9 × 2) = √9 × √2 = 3√2

OpenStudy (anonymous):

ok thank you :) <3

OpenStudy (anonymous):

Your welcome!!!